OpenStudy (anonymous):

For an experiment in science class, students roll a number cube with faces numbered 1 to 8, then draw a card from deck of 3 cards - the ace of hearts, queen of spades, and eight of clubs, and then finally flip a coin. How many possible outcomes are there for this experiment ? PLZZ HELP !!!!

OpenStudy (welshfella):

How many possible outcomes are there when you roll the cube of 8 faces?

OpenStudy (mathstudent55):

Use the counting principle.

OpenStudy (anonymous):

@mathstudent55 what is the counting principle ????

OpenStudy (anonymous):

@welshfella im not sure

OpenStudy (welshfella):

there are 8 numbers right? so if you roll the dice any one of 1 to 8 will be obtaines So that is how many outcomes?

OpenStudy (anonymous):

8 @welshfella

OpenStudy (welshfella):

right so for every 1 of the eight numbers you will get any 1 of the 3 cards - that's a possible number of outcomes of 8 * 3 = 24. (this is the counting principle)

OpenStudy (welshfella):

Flipping a coin has 2 outcomes Heads and tails so possible outcomes for the whole experiment = 24 * what?

OpenStudy (anonymous):

Um 24*8 ??

OpenStudy (welshfella):

no - we are dealing with the coin now - 2 outcomes

OpenStudy (welshfella):

for everyone of the 24 outcomes we can get a head or a tail

OpenStudy (anonymous):

24*2

OpenStudy (welshfella):

so it 24 * ?

OpenStudy (welshfella):

right

OpenStudy (anonymous):

Which is 48

OpenStudy (welshfella):

its 8*3*2 = 48

OpenStudy (welshfella):

yea

OpenStudy (anonymous):

That's the final answer ?

OpenStudy (anonymous):

@welshfella

OpenStudy (mathstudent55):

The counting principle is find how many ways each thing can happen and multiply them all together.

OpenStudy (mathstudent55):

cube: 8 outcomes cards: 3 outcomes coin: 2 outcomes total outcomes: 8 * 3 * 2 = 48 You are correct.

OpenStudy (anonymous):

Thanks @mathstudent55

OpenStudy (mathstudent55):

You're welcome.